Unit 28 Homework Assignment  Physics141-Summer 2009

Unit 28 homework after session 1

• Read chapter 34 in the textbook Understanding Physics up to section 34-7.

• Work textbook problems 34-3, 34-11, 34-12 and 34-14. Unit 28 homework after session 2

• Read the rest of chapter 34 in the textbook ”Understanding Physics”.

• Work textbook problems 34-27, 34-39, and 34-55.

All Unit 28 homework problems and the unit 28 activity guide are due on the first day of Unit 29.